OpenStudy (anonymous):

multiply and simplify: [(x+3)+5y][(x+3)-5y]

OpenStudy (anonymous):

\[(a+b)(a-b)=a^2-b^2\] with \(a=(x+3)\) and \(b=5y\)

OpenStudy (anonymous):

(x+3)^2 -25y?

OpenStudy (anonymous):

No what is the square of 5y??

OpenStudy (anonymous):

5y^2?

OpenStudy (anonymous):

No again check your answer...

OpenStudy (anonymous):